# Rings of non-commutative functions and their fields of fractions

### v1

@@ 4 Algebraic properties of rings of analytic NC functions
Let $F$ and $G$ be as above, and let $r_{0}$ be the number obtained in the preceding lemma. Given the strictly increasing sequence $(t_{m})_{m=1}^{\infty}$, $0<t_{m}<R$, $t_{m}\uparrow R$, fix $m\in\mathbb{N}$, such that $t_{m}>r_{0}$. Let $S\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m+1}))$ and $S^{\prime}\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m}))$ be such that
$FS=\begin{pmatrix}\tilde{F}&0\end{pmatrix},\ FS^{\prime}=\begin{pmatrix}\tilde{F}^{\prime}&0\end{pmatrix},\qquad\mbox{and}\qquad S^{-1}G=\begin{pmatrix}0\\ \tilde{G}\end{pmatrix},\ (S^{\prime})^{-1}G=\begin{pmatrix}0\\ \tilde{G}^{\prime}\end{pmatrix}$,
in the sense that the first two equations hold for all evaluations in $t_{m+1}\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ and the second two hold for all evaluations in $t_{m}\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, $\widetilde{F}\in\mathbb{A}_{d}(t_{m+1})\otimes\mathbb{C}^{1\times n}$, $\widetilde{G}\in\mathbb{A}_{d}(t_{m+1})\mathbb{C}^{n}$, $\widetilde{F}^{\prime}\in\mathbb{A}_{d}(t_{m})\otimes\mathbb{C}^{1\times n}$ and $\widetilde{G}^{\prime}\in\mathbb{A}_{d}(t_{m})\otimes\mathbb{C}^{n}$.
@@
We conclude, from Lemma 4.1, that there exists a lower triangular $T^{\prime}\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m}))$, such that $S^{\prime}=\pi_{m+1,m}(S)T^{\prime}$. Now for every $\varepsilon>0$, we can choose $T\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m+1}))$ lower triangular, such that $\|\pi_{m+1,m}(T)-T^{\prime}\|<\frac{\varepsilon}{\|S\|}$. Indeed,
@@
Since both $S^{\prime}$ and $\pi_{m+1,m}(S)$ trivialize our relation $FG=0$ on the row-ball of radius $t_{m}$, Lemma 4.1 implies that there is a block lower triangular $\widetilde{T}^{\prime}$ so that $\pi_{m+1,m}(S)=S^{\prime}\widetilde{T}^{\prime}$. It follows that $\widetilde{T}^{\prime}=T^{{}^{\prime};-1}$, so that $T^{{}^{\prime};-1}$ is also block lower triangular (with the same sized blocks). If
@@
Note that the images of $\mathbb{A}_{d}(t_{m+1})^{n\times n}$ and $\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m+1}))$ in $\mathbb{A}_{d}(t_{m})^{n\times n}$ and $\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m}))$, respectively, are dense. Indeed, this can be see from the fact any element of the unital operator algebra, $\mathbb{A}_{d}(t_{m})^{n\times n}$, is the supremum–norm limit of elements of $\mathbb{C}\langle\mathfrak{z}\rangle^{n\times n}$ and the group of invertible elements in a unital Banach algebra is an open set. Hence, there exist matrices $A,B,C$ over $\mathbb{A}_{d}(t_{m+1})$ of appropriate sizes so that $A,C$ are invertible and so that their images under the restriction map, $\pi_{m+1,m}$ are arbitrarily close to $A^{\prime},B^{\prime},C^{\prime}$, respectively. That is, for any $\epsilon>0$, we can choose
@@
In other words, we can replace $S$ with a different trivializing matrix, such that its image will be close to the solution $T$ within the desired arbitrary precision.
@@
    For all $m\in\mathbb{N}$, $F|_{m}S_{m}=(\tilde{F}_{m},0_{\ell})$ with $\tilde{F}_{m}$ injective, and where $F|_{m}$ denotes the restriction of $F$ to the row-ball of radius $t_{m}$,
@@
For each $m>m_{0}$, by Lemma 4.2, we can find $T_{m}\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m}))$, such that $\pi_{m}(F)T_{m}=(\tilde{F}_{m},0_{\ell})$, with $\tilde{F}^{\prime}_{m}$ injective. Now for $m=1$, we set $S_{m}=T_{m_{0}+1}$ and let $\varepsilon_{1}=\min\{1,r_{m_{0}+1}(S_{1})/2\}$. Now assume that $S_{1},\ldots,S_{m}$ and $\varepsilon_{1},\ldots,\varepsilon_{m}$ have been defined. By the argument above, an $S_{m+1}\in\operatorname{GL}_{n}(\mathbb{A}_{d}(m_{0}+m+1))$ exists, such that $\|\pi^{(n)}_{m_{0}+m+1,m_{0}+m}(S_{m+1})-S_{m}\|<\varepsilon_{m}$ and $\pi_{m_{0}+m+1}(F)S_{m+1}=(\tilde{F}_{m},0_{\ell})$, with $\tilde{F}_{m+1}$ injective. Now we set $\varepsilon_{m+1}=\frac{1}{2^{m}}\min\{1,r_{m_{0}+1}(S_{1})/2,\ldots,r_{m_{0}+m+1}(S_{m+1})/2\}$. The first and second properties are satisfied immediately from the construction. To check the last property, let us fix an arbitrary $m\in\mathbb{N}$. By construction, $\varepsilon_{m+k}$ is a minimum taken over a finite set of numbers, one of which is $r_{m_{0}+m}(S_{m})/2$ divided by $2^{m+k-1}$.
Having these sequences at hand, we fix $m\in\mathbb{N}$ and set $S_{m,0}^{\prime}=S_{m}$. For every $k\in\mathbb{N}$, we set $S_{m,k}^{\prime}=\pi_{m+m_{0}+k,m+m_{0}}^{(n)}(S_{m+k})$. Since every $\pi_{m_{0}+m+1,m_{0}+m}$ is a complete contraction, then so is $\pi_{m_{0}+m+k,m_{0}+m}$. We note now that
@@
In particular, the Cauchy sequence stays in a closed ball of radius $\frac{r_{m+m_{0}}(S_{m})}{2^{m}}$ around $S_{m}$ and, thus, its limit is invertible. Let $S_{m}^{\prime}$ be the limit. We now compute,
@@
$=$ $\lim_{k\to\infty}(S_{m,k+1}^{\prime})=S_{m}^{\prime}.$
Therefore, the inverse sequence, $S_{m}^{\prime}$ defines an element $S^{\prime}\in\mathscr{O}_{d}(R)^{n\times n}$. Moreover, since every $S_{m}^{\prime}$ is invertible, $\pi_{m+m_{0}+1,m+m_{0}}^{(n)}(S_{m+m_{0}+1}^{{}^{\prime};-1})=S_{m}^{{}^{\prime}:-1}$ and we conclude that $S^{\prime}\in\operatorname{GL}_{n}(\mathscr{O}_{d}(R))$. Lastly,
$FS_{m}^{\prime}=\lim_{k\to\infty}FS_{m,k}^{\prime}=\lim_{k\to\infty}\begin{pmatrix}\underbrace{\star}_{n-\ell}&0_{\ell}\end{pmatrix}$
@@
Let $F\in\mathscr{O}_{d}^{1\times n}(R)$, $R\in(0,+\infty]$. Then the map $F\colon\mathscr{O}_{d}(R)\otimes\mathbb{C}^{n}\to\mathscr{O}_{d}(R)$ is injective if and only if $F(rL)$ is injective for sufficiently large $0<r<R$.
@@
Clearly, if $F$ is not injective, then $F(rL)$ is not injective for every $0<r<R$. Now assume that $F(rL)$ is not injective for every $0<r<R$. Let $\ell(r)=\mathrm{dim}\,\mathscr{W}_{\mathrm{Ker}\,F(rL)}$. As in the proof of Lemma 4.2, the function $\ell(r)$ is monotonically decreasing, integer-valued, and bounded below by $1$ by our assumption. Therefore, there exists $0<r_{0}<R$, so that for $0<r_{0}\leq r<R$, $\ell(r)=:\ell\in\mathbb{N}$ is constant. Therefore, for $0<r_{0}\leq r<s<R$, the restriction map from $\mathrm{Ker}\,F(sL)$ to $\mathrm{Ker}\,F(rL)$ has dense range. In other other words, for every $\varepsilon>0$ and every $G_{r}\in\mathrm{Ker}\,F(rL)$, there exists $H\in\mathrm{Ker}\,F(sL)$, such that $\|H(\tfrac{r}{s}L)1-G_{r}\|<\varepsilon$. Now we argue, as in the proof of Theorem 4.3. Given the strictly increasing sequence $(t_{m})_{m\in\mathbb{N}}$ so that $0<t_{m}<R$ and $t_{m}\uparrow\mathbb{R}$, let $m_{0}\in\mathbb{N}$ be the minimal value so that $t_{m_{0}}\geq r_{0}$. Then for any $m>m_{0}$, we will inductively build sequences $\varepsilon_{m}>0$ and $G_{m}\in\mathrm{Ker}\,F(mL)$ with the properties:
@@
If we show that $J$ is closed, then by Lemma 3.6, we can deduce that $J=I(\mathscr{Z}(J))$. Since $J$ is finitely-generated and $\mathscr{O}_{d}(R)$ is a semifir, $J$ is free as a right $\mathscr{O}_{d}(R)$-module. Namely, there exists $n\in\mathbb{N}$ and $F\in\mathscr{O}_{d}(R)^{1\times n}$, such that $F\cdot(\mathscr{O}_{d}(R)\otimes\mathbb{C}^{n})=J$ and $F$ is injective. Assume that $FG_{n}\to h\in\mathscr{O}_{d}(R)$. Since $F$ is injective, by the preceding lemma, there exists $r_{0}>0$, such that for every $0<r_{0}\leq r<R$, $F(rL)$ is injective. By increasing $r_{0}$, if necessary, we may apply Lemma 2.5. By the definition of the topology on $\mathscr{O}_{d}(R)$, we have that $F(rL)G_{n}(rL)\to h(rL)$ in $\mathbb{A}_{d}$ for any $0<r<R$. Let $0<r_{0}\leq r<s<s^{\prime}<R$. Write $F(s^{\prime}L)=VT$, the inner–outer decomposition. Since $h\in\mathscr{O}_{d}(R)$, $h(s^{\prime}L)$ belongs to $\mathbb{A}_{d}$ for any $0<s^{\prime}<R$. In particular, since $VTG_{n}(s^{\prime}L)1\to h(s^{\prime}L)1$ in $\mathbb{H}^{2}_{d}$, $TG_{n}(s^{\prime}L)\rightarrow V^{*}h(s^{\prime}L)1$ in $\mathbb{H}^{2}_{d}$, so that $T(\frac{s}{s^{\prime}}L)G_{n}(sL)$ converges in $\mathbb{A}_{d}$ for any $0<r_{0}\leq r<s<s^{\prime}<R$. By Lemma 2.5, $T(\tfrac{r}{s^{\prime}}L)$ is square and invertible. Hence, $G_{n}(rL)$ converges in $\mathbb{A}_{d}$. In particular, for every $0<r_{0}\leq r<R$, $G_{n}(rL)$ is bounded. Therefore, by Theorem 3.3, $G_{n}$ has a subsequence that converges in $\mathscr{O}_{d}(R)\otimes\mathbb{C}^{n}$ to some $G$. By the continuity of multiplication $h=FG\in J$. Hence, $J$ is closed. ∎
@@
The stabilization of the wandering dimension used both in the proof of Lemma 2.5 and Lemma 4.4 can be viewed as analogous to the algebraic Mittag–Leffler condition on inverse limits [49, Section 0594].

### v2

@@ 4 Algebraic properties of rings of analytic NC functions
Let $F$ and $G$ be as above, and let $r_{0}$ be the number obtained in the preceding lemma so that for any $r\geq r_{0}$, the wandering dimension of $\mathrm{Ker}\,F(rL)$ is constant and equal to $\ell\in\mathbb{N}$, $\ell<n$. Given the strictly increasing sequence $(t_{m})_{m=1}^{\infty}$, $0<t_{m}<R$, $t_{m}\uparrow R$, fix $m\in\mathbb{N}$, such that $t_{m}>r_{0}$. By Lemma 4.2, there exist $S\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m+1}))$ and $S^{\prime}\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m}))$ be such that
$FS=\begin{pmatrix}\tilde{F}&0\end{pmatrix},\ S^{-1}G=\begin{pmatrix}0\\ \tilde{G}\end{pmatrix},\qquad\mbox{and}\qquad FS^{\prime}=\begin{pmatrix}\tilde{F}^{\prime}&0\end{pmatrix},\ (S^{\prime})^{-1}G=\begin{pmatrix}0\\ \tilde{G}^{\prime}\end{pmatrix}$,
in the sense that the first two equations hold for all evaluations in $t_{m+1}\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ and the second two hold for all evaluations in $t_{m}\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, $\widetilde{F}\in\mathbb{A}_{d}(t_{m+1})\otimes\mathbb{C}^{1\times(n-\ell)}$, $\widetilde{G}\in\mathbb{A}_{d}(t_{m+1})\otimes\mathbb{C}^{\ell}$, $\widetilde{F}^{\prime}\in\mathbb{A}_{d}(t_{m})\otimes\mathbb{C}^{1\times(n-\ell)}$ and $\widetilde{G}^{\prime}\in\mathbb{A}_{d}(t_{m})\otimes\mathbb{C}^{\ell}$. Moreover, by the aforementioned lemma, $\tilde{F}$ and $\tilde{F^{\prime}}$ are both injective.
@@
We conclude, from Lemma 4.1, that for any $m\geq m_{0}$, there exists a block lower triangular $T^{\prime}=\left(\begin{smallmatrix}A^{\prime}&0\\ B^{\prime}&C^{\prime}\end{smallmatrix}\right)\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m}))$, with $A^{\prime}\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{(n-\ell)\times(n-\ell)}$, $C^{\prime}\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{\ell\times\ell}$, such that $S^{\prime}=\pi_{m+1,m}(S)T^{\prime}$. For every $\varepsilon>0$, we can choose a block lower triangular $T\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m+1}))$ so that $\|\pi_{m+1,m}(T)-T^{\prime}\|<\frac{\varepsilon}{\|S\|}$. Indeed,
@@
Since both $S^{\prime}$ and $\pi_{m+1,m}(S)$ trivialize our relation $F\cdot G=0$ on the row-ball of radius $t_{m}$, Lemma 4.1 implies that there is a block lower triangular $\widetilde{T}^{\prime}$ so that $\pi_{m+1,m}(S)=S^{\prime}\widetilde{T}^{\prime}$. It follows that $\widetilde{T}^{\prime}=T^{{}^{\prime};-1}$, so that $T^{{}^{\prime};-1}$ is also block lower triangular (with the same sized blocks). If
@@
Note that the images of $\mathbb{A}_{d}(t_{m+1})^{n\times n}$ and $\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m+1}))$ in $\mathbb{A}_{d}(t_{m})^{n\times n}$ and $\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m}))$, respectively, are dense. The latter statement was proved in Lemma 3.6. Hence, there exist matrices $A,B,C$ over $\mathbb{A}_{d}(t_{m+1})$ of appropriate sizes so that $A,C$ are invertible and so that their images under the restriction map, $\pi_{m+1,m}$ are arbitrarily close to $A^{\prime},B^{\prime},C^{\prime}$, respectively. That is, for any $\epsilon>0$, we can choose
@@
In other words, we can replace $S$ with a different trivializing matrix, such that its image will be close to the solution $T$ within the desired arbitrary precision in the row-ball of radius $t_{m}$.
@@
    For all $m\in\mathbb{N}$, $F_{m}S_{m}=(\tilde{F}_{m},0_{\ell})$ with $\tilde{F}_{m}$ injective, and where $F_{m}$ denotes the restriction of $F$ to the row-ball of radius $t_{m}$,
@@
For each $m>m_{0}$, by Lemma 4.2, we can find $T_{m}\in\operatorname{GL}_{n}(\mathbb{A}_{d}(t_{m}))$, such that $\pi_{m}(F)T_{m}=(\tilde{F}_{m},0_{\ell})$, with $\tilde{F}^{\prime}_{m}$ injective. For $m=1$, we set $S_{m}=T_{m_{0}+1}$ and let $\varepsilon_{1}=\min\{1,r_{m_{0}+1}(S_{1})/2\}$. Assume that $S_{1},\cdots,S_{m}$ and $\varepsilon_{1},\cdots,\varepsilon_{m}$ have been defined. By the argument above, an $S_{m+1}\in\operatorname{GL}_{n}(\mathbb{A}_{d}(m_{0}+m+1))$ exists, such that $\|\pi^{(n)}_{m_{0}+m+1,m_{0}+m}(S_{m+1})-S_{m}\|<\varepsilon_{m}$ and $\pi_{m_{0}+m+1}(F)S_{m+1}=(\tilde{F}_{m},0_{\ell})$, with $\tilde{F}_{m+1}$ injective. We now set $\varepsilon_{m+1}=\frac{1}{2^{m}}\min\{1,r_{m_{0}+1}(S_{1})/2,\cdots,r_{m_{0}+m+1}(S_{m+1})/2\}$. The first and second properties are satisfied immediately from the construction. To check the last property, let us fix an arbitrary $m\in\mathbb{N}$. By construction, $\varepsilon_{m+k}$ is a minimum taken over a finite set of numbers, one of which is $r_{m_{0}+m}(S_{m})/2$ divided by $2^{m+k-1}$.
Having these sequences at hand, we fix $m\in\mathbb{N}$ and set $S_{m,0}^{\prime}:=S_{m}$. For every $k\in\mathbb{N}$, we further set $S_{m,k}^{\prime}:=\pi_{m+m_{0}+k,m+m_{0}}^{(n)}(S_{m+k})$. Since every $\pi_{m_{0}+m+1,m_{0}+m}$ is a complete contraction, so is $\pi_{m_{0}+m+k,m_{0}+m}$. We note now that
@@
In particular, the Cauchy sequence stays in a closed ball of radius $\frac{r_{m+m_{0}}(S_{m})}{2^{m}}$ around $S_{m}$ and hence its limit, $S_{m}^{\prime}$, is invertible. We now compute,
@@
$=$ $\lim_{k\to\infty}S_{m,k+1}^{\prime}=S_{m}^{\prime}.$
This shows that the sequence, $S_{m}^{\prime}$, is an inverse limit sequence and therefore defines an element, $S^{\prime}\in\mathscr{O}_{d}(R)^{n\times n}$. Moreover, since every $S_{m}^{\prime}$ is invertible, $\pi_{m+m_{0}+1,m+m_{0}}^{(n)}(S_{m+m_{0}+1}^{{}^{\prime};-1})=S_{m}^{{}^{\prime}:-1}$ and we conclude that $S^{\prime}\in\operatorname{GL}_{n}(\mathscr{O}_{d}(R))$. Lastly,
$FS_{m}^{\prime}=\lim_{k\to\infty}FS_{m,k}^{\prime}=\lim_{k\to\infty}(\underbrace{\star}_{n-\ell}\ 0_{\ell})$
@@
Let $F\in\mathscr{O}_{d}^{1\times n}(R)$, $R\in(0,+\infty]$. Then, the map $F\colon\mathscr{O}_{d}(R)\otimes\mathbb{C}^{n}\to\mathscr{O}_{d}(R)$ is injective if and only if $F(rL)$ is injective for sufficiently large $0<r<R$.
@@
Clearly, if $F$ is not injective, then $F(rL)$ is not injective for every $0<r<R$. Now assume that $F(rL)$ is not injective for every $0<r<R$. Let $\ell(r)=\mathrm{dim}\,\mathscr{W}_{\mathrm{Ker}\,F(rL)}$ be the wandering dimension of $\mathrm{Ker}\,F(rL)$. As in the proof of Lemma 4.2, the function $\ell(r)$ is monotonically decreasing, integer-valued, and bounded below by $1$ by our assumption. Therefore, there exists $0<r_{0}<R$, so that for $0<r_{0}\leq r<R$, $\ell(r)=:\ell\in\mathbb{N}$ is constant. Therefore, for $0<r_{0}\leq r<s<R$, the restriction map from $\mathrm{Ker}\,F(sL)$ to $\mathrm{Ker}\,F(rL)$ has dense range. In other other words, for every $\varepsilon>0$ and every $G_{r}\in\mathrm{Ker}\,F(rL)$, there exists $H\in\mathrm{Ker}\,F(sL)$, such that $\|H(\tfrac{r}{s}L)1-G_{r}\|<\varepsilon$. Now we argue as in the proof of Theorem 4.3. Given the strictly increasing sequence $(t_{m})_{m\in\mathbb{N}}$ so that $0<t_{m}<R$ and $t_{m}\uparrow R$, let $m_{0}\in\mathbb{N}$ be the minimal value so that $t_{m_{0}}\geq r_{0}$. Then for any $m>m_{0}$, we inductively build sequences $\varepsilon_{m}>0$ and $G_{m}\in\mathrm{Ker}\,F(mL)$ with the properties:
@@
If we show that $J$ is closed, then by Lemma 3.7, we can deduce that $J=I(\mathscr{Z}(J))$. Since $J$ is finitely-generated and $\mathscr{O}_{d}(R)$ is a semifir, $J$ is free as a right $\mathscr{O}_{d}(R)$-module. Namely, there exists $n\in\mathbb{N}$ and $F\in\mathscr{O}_{d}(R)^{1\times n}$, such that $F\cdot(\mathscr{O}_{d}(R)\otimes\mathbb{C}^{n})=J$ and $F$ is injective. Assume that $FG_{n}\to h\in\mathscr{O}_{d}(R)$. Since $F$ is injective, by the preceding lemma, there exists $r_{0}>0$, such that for every $0<r_{0}\leq r<R$, $F(rL)$ is injective. By increasing $r_{0}$, if necessary, we may apply Lemma 2.5. By the definition of the topology on $\mathscr{O}_{d}(R)$, we have that $F(rL)G_{n}(rL)\to h(rL)$ in $\mathbb{A}_{d}$ for any $0<r<R$. Let $0<r_{0}\leq r<s<s^{\prime}<R$. Write $F(s^{\prime}L)=VT$, the inner–outer decomposition. Since $h\in\mathscr{O}_{d}(R)$, $h(s^{\prime}L)$ belongs to $\mathbb{A}_{d}$ for any $0<s^{\prime}<R$. In particular, since $V(L)T(L)G_{n}(s^{\prime}L)1\to h(s^{\prime}L)1$ in $\mathbb{H}^{2}_{d}$, $T(L)G_{n}(s^{\prime}L)\rightarrow V(L)^{*}h(s^{\prime}L)1$ in $\mathbb{H}^{2}_{d}$, so that $T(\frac{s}{s^{\prime}}L)G_{n}(sL)$ converges in $\mathbb{A}_{d}$ for any $0<r_{0}\leq r<s<s^{\prime}<R$. By Lemma 2.5, $T(\tfrac{r}{s^{\prime}}L)$ is square and invertible. Hence, $G_{n}(rL)$ converges in $\mathbb{A}_{d}$. In particular, for every $0<r_{0}\leq r<R$, $G_{n}(rL)$ is bounded. Therefore, by Theorem 3.3, $G_{n}$ has a subsequence that converges in $\mathscr{O}_{d}(R)\otimes\mathbb{C}^{n}$ to some $G$. By the continuity of multiplication $h=F\cdot G\in J$. Hence, $J$ is closed. ∎
@@
The stabilization of the wandering dimension used both in the proof of Lemma 2.5 and Lemma 4.4 can be viewed as analogous to the algebraic Mittag–Leffler condition on inverse limits [51, Section 0594].